### v1

# Materials and spin characteristics of amino-terminated nanodiamonds embedded with nitrogen-vacancy color centers

Abstract
Understanding and controlling fluorescent nanodiamonds (FNDs) with optically read qubits is a key focus of research, as they show high potential for detecting electric and magnetic fields, temperature, and other physico-chemical quantities at the nanoscale, which are highly sought after in chemistry and biology. Proper functionalization of FNDs is required for their application as probes in chemical and biological processes. However, modifying the surface of FNDs can affect the properties of qubit sensors. In this work, we thoroughly study the fundamental properties of embedded nitrogen-vacancy (NV) color centers as a function of FND size and surface termination. The FNDs were produced by milling high-pressure, high-temperature diamonds, and NV centers were introduced via electron beam irradiation and annealing. In particular, the initial FNDs, covered with various oxygen groups, were homogenized through a reduction process to predominantly cover them with alcohol (--OH) groups as reference FNDs. Additionally, a Hofmann degradation process was applied to terminate FNDs with amine (--NH2) groups, enabling direct linkage of proteins and other biomolecules to the FNDs. We monitored the charge state stability upon illumination, the zero-field splitting parameters, and the longitudinal spin-relaxation time of the NV centers in these FNDs. Our findings indicate that charge state stabilization of the NV centers was achieved in both --OH- and --NH2-terminated FNDs beyond a certain FND size. Furthermore, we demonstrate that --NH2-terminated FNDs embedding NV centers exhibit a $T_{1}$ longitudinal relaxation time of approximately 25 $\mu$s, independent of FND size within the studied range of 10 nm to 140 nm. This relaxation time is comparable to that of larger-sized reference FNDs. We attribute the unique size-independent $T_{1}$ behavior of amino-terminated FNDs to the complex interplay of surface chemistry during the Hofmann degradation process. Our amino-terminated FNDs have promising potential for functionalization in biochemical targeting and simultaneous application as quantum sensor probes.
diamond nanoparticle, fluorescent nanoparticle, defects in diamond, fluorescence lifetime, FTIR spectroscopy, --OH termination, --NH2 termination, XPS, EPR
I Introduction
Nitrogen-vacancy (NV) centers engineered into nanodiamonds are leading contenders to realize quantum sensor probes for chemistry and biology Kucsko et al. (2013); Rendler et al. (2017); Fujisaku et al. (2019); Perona Martínez et al. (2020); Webb et al. (2020); Barton et al. (2020); Zhang et al. (2021); Wu and Weil (2022); Sigaeva et al. (2022); Liu et al. (2023); Hart and Knowles (2023); Belser et al. (2023); Gu et al. (2023); Oshimi et al. (2024); Lu et al. (2024). The NV center is a photostable fluorescent defect or color center in diamond so nanodiamonds embedded with NV centers are often called fluorescent nanodiamonds (FNDs) Yu et al. (2005); Hsiao et al. (2016). Besides the optical activity of the NV center it has other unique magneto-optical properties that can be employed to realize quantum sensors from FNDs. In particular, its electronic spin state can be polarized and read-out optically where the relaxation time of the electron spin of the NV centers in FNDs can be used to detect the presence of electron spins external to FNDs, e.g., radicals or other paramagnetic species ejected in cells. Biological application of FND quantum sensors requires a tight control of the surface of nanodiamonds, in order to simultaneously maintain the exquisite properties of the NV quantum sensor and functionalize the FNDs to study the desired biological molecules. To depict the complexity of this issue, we provide a detailed description about FNDs embedded with NV centers and the motivation to study amino-terminated FNDs in the following paragraphs.
The NV center contains a nitrogen substituting carbon adjacent to a vacancy in diamond where the negatively charged NV defect has $S=1$ paramagnetic ground state and its spin-selective fluorescence makes it possible to optically spinpolarize and read out its spin state, by i.e., optically detected magnetic resonance (ODMR) of the electron spin where the $m_{S}=0$ and $m_{S}=\pm 1$ spin levels split (2.87 GHz) at zero magnetic field due to the axial $C_{3v}$ symmetry of the NV center Doherty et al. (2013); Gali (2019). The negative charge state is provided by the substitutional nitrogen donors near the NV center in the diamond. The $S=1$ spin state has few milliseconds (ms) longitudinal spin relaxation time ($T_{1}$ time) at room temperature for deeply buried NV centers in diamond Rosskopf et al. (2014); Alsid et al. (2019) where the room-temperature ODMR readout contrast for single NV center can reach 30% when the NV center is photoexcited to the phonon sideband (for excitation one can apply e.g., $514\dots 560$ nm wavelengths of the laser) and the emitted photons are observed in the phonon sideband, e.g., in the region of 650 and 900 nm wavelengths. The Hahn-echo Hahn (1950) $T_{2}$ coherence time of the NV electron spin can principally be similarly long as the $T_{1}$ time Balasubramanian et al. (2009); Cambria et al. (2023a). The spin dephasing or free-induction decay $T_{2}^{*}$ time is much shorter than $T_{2}$ time in diamonds with natural abundance of carbon isotopes that is often caused by the fluctuating magnetic fields of 13C $I=1/2$ nuclear spins (typically, $T_{2}^{*}\approx 2$ $\mu$s at zero constant magnetic field Zhao et al. (2012)). As a consequence, the coherence and spin dephasing times could be elongated by significant reduction of the 13C isotope density in diamond at zero or near-zero magnetic fields Balasubramanian et al. (2009).
These electron spin-related parameters are sensitive for magnetic, electric and strain fields (see Ref. Barry et al., 2020 and references therein) as well as temperature Acosta et al. (2010); Jarmola et al. (2012); Lin et al. (2021); Cambria et al. (2023a, b), thus NV center can be employed as quantum sensor at the nanoscale where optical methods are combined with magnetic resonance techniques to carry out ODMR detection of NV centers. Room temperature operation and spin characteristics of the NV center make it a very attractive platform to realize quantum sensor probes for biology. To this end, NV centers may be engineered into small nanodiamonds where NV centers per se reside close to the surface to sense the environment Tisler et al. (2009); Rondin et al. (2010); Bradac et al. (2010); Terada et al. (2019), and these nanodiamonds can be delivered into living organism that can be tracked, even at individual level, by optical means Mohan et al. (2010); McGuinness et al. (2011); Igarashi et al. (2012); Hegyi and Yablonovitch (2013); Yanagi et al. (2021).
Although, NV color center is considered to be photostable defect but blinking of NV center is observed, e.g., in 5-nm sized FNDs Bradac et al. (2010). This is associated with the charge switching of the NV defect from negative, NV(--) to neutral, NV(00). NV(00) yields distinct photoluminescence (PL) spectrum from that of NV(--Mita (1996), so the charge switching can be monitored in PL spectroscopy even for individual NV defect (e.g., Ref. Neethirajan et al., 2023). This effect is associated with the downward band bending caused by the negatively charged species at the diamond surface which will stabilize the neutral charge state of the NV defect or even its positive charge state in extreme cases beneath the diamond surface Pfender et al. (2017); Broadway et al. (2018). Recent first principles calculations imply that illumination employed to drive near-surface NV centers leads to the charge switching of the vacancy clusters – generated in proximity to the NV defects during the preparation of the NV defects by irradiation and annealing of diamond Deák et al. (2014); Fávaro de Oliveira et al. (2017) – which plays a crucial role in the charge switching process of the near-surface NV defects Neethirajan et al. (2023); Freire-Moschovitis et al. (2023).
Besides the photostability of NV center, the spin properties of near-surface NV centers are also affected. The presence of nuclear spins, e.g., 1H $I=1/2$ spins naturally shorten the $T_{2}^{*}$ time of near-surface NV centers via dipole-dipole hyperfine coupling between the electron spin of the NV center and the nuclear spins at the surface. In larger nanodiamonds, where NV centers reside farther from the surface, the proximate nitrogen donor spins inside nanodiamonds – the so-called P1 electron spin resonance (ESR) centers Loubser and Wyk (1978) – reduce the $T_{2}^{*}$ time of the NV center via electron spin - electron spin interaction Knowles et al. (2014). The charge and electron spin fluctuations between the nearby vacancy clusters around the NV center shorten its $T_{1}$ time Freire-Moschovitis et al. (2023) when compared to that of deeply buried NV centers, and the rate of charge fluctuation near the NV center so the $T_{1}$ time of the NV center can be influenced by engineering the chemical species at the diamond surface Freire-Moschovitis et al. (2023).
Fluorescent nanodiamonds can be prepared by various techniques Shenderova et al. (2019). The resulting shape of nanodiamonds is often irregular and far from spherical. This behavior stands out for nanodiamonds that are prepared by top-down techniques, e.g. milling from large diamond particles. The size of these nanodiamonds are usually measured as the largest distance between the facets of the particle. Because of the irregular shape of FNDs, the distance of the NV defect from the nanodiamond surface cannot be well estimated. As a consequence, picking up single NV centers from the same size of oxygenated FNDs yields an order of magnitude variance in the observed $T_{1}$ times, for the considered FND sizes of $\approx 10-70$ nm; on the other hand, the observed longitudinal spin relaxation rate falls $10^{3}$-$10^{4}$ Hz ($T_{1}\approx 0.1$-$1$ ms) for 60-70 nm sized FNDs whereas it typically goes as $4\times 10^{3}$-$2\times 10^{5}$ Hz ($T_{1}\approx 5$-$250$ $\mu$s) for smaller FNDs Tetienne et al. (2013). The observed temperature dependence of the $T_{1}$ time in small FNDs showed a moderate variance; it only became 2.5$\approx$ 2.5$\times$ 2.5 × longer at 4 K than that at room temperature de Guillebon et al. (2020). The origin of this effect is not clear but it is unlikely related to spin-phonon relaxation of the NV center because that changes the $T_{1}$ value with few orders of magnitudes in this range of temperatures as observed in bulk diamond Jarmola et al. (2012); Cambria et al. (2023a).
The NV center in FNDs can be employed as efficient quantum sensors when the favorable optical and spin properties in bulk diamond can be preserved in FNDs. The above mentioned studies imply that reducing the 13C $I=1/2$ nuclear spin density, optimizing the nitrogen concentration for maintaining NV(--) defects but minimizing the nitrogen donor spins’ noise at the same time, and optimizing the surface of diamond Kaviani et al. (2014); Sangtawesin et al. (2019) are feasible routes to this end. Indeed, relatively large nanodiamonds ($\approx 100$ nm in size) have been recently produced by milling polycrystalline 12C enriched chemical vapor deposited diamonds March et al. (2023) where the estimated concentration of the P1 ESR centers was about $0.15\pm 0.02$ parts per million (ppm). In these FNDs, the observed $T_{1}$, $T_{2}$, and $T_{2}^{*}$ values approached those observed in bulk diamond with similar concentration of P1 centers March et al. (2023). A similar strategy has been applied to high-pressure high-temperature (HPHT) 12C enriched FNDs with $\approx 200$ nm of size Oshimi et al. (2024) where the estimated nitrogen donor concentration was $\approx 30$-$60$ ppm. As a consequence, the average $T_{1}=0.68$ ms and $T_{2}=3.2$ $\mu$s times of the observed NV centers in these FNDs demonstrate a significant improvement but they are shorter because of the higher concentration of the P1 centers than those in Ref. March et al., 2023. For smaller FNDs, the surface species may also pose a problem for the quality of NV centers, in particular, for the $T_{1}$ time Rondin et al. (2010); Ryan et al. (2018); Freire-Moschovitis et al. (2023); Zvi et al. (2023). Another approach to create high-quality diamond embedded with NV center is the CVD growth of diamond with nitrogen doping where nano-roughening can be used to nucleate diamond nanoparticles Prooth et al. (2023). The resulting FNDs are relatively large (few hundreds of nanometers) but relatively small ones ($\sim 60$ nm) can be also produced with embedding NV centers with long $T_{1}$ times (mean value is 800 $\mu$s) Prooth et al. (2023). In these FNDs the concentration of P1 centers should be relatively low that could lead to long $T_{1}$ times of NV centers.
Along this line of research, Type-1b HPHT diamonds (with non-specified concentration of P1 centers) were used in a previous study Ryan et al. (2018) where the $T_{1}$ time of 10 single NV centers was monitored in oxygenated and borane-reduced FNDs (50 nm in size). It was found that the borane reduction process Krüger et al. (2006); Krueger et al. (2008) causes a loss of C=O and an associated production of C–O groups and produces C–H groups by reacting with sp2 carbon at the surface of nanodiamonds Ryan et al. (2018). The reduction of the oxidized surface is suitable for both elemental hydrogen and electrochemical methods. The results on the $T_{1}$ time was not entirely conclusive as the $T_{1}$ time was reduced for two NV centers but it increased for the other NV centers with various amounts where the average enhancement in $T_{1}$ time including all the 10 NV centers in the sampling was a factor of two Ryan et al. (2018).
Motivated by theoretical studies suggesting the mixture of ether, alcohol (--OH) and --H groups on the (100) diamond surface could be optimal for NV quantum sensors Kaviani et al. (2014), oxygenated HPHT FNDs were treated by non-thermal plasma with H2 working gas in a recent study Gulka et al. (2024) which reduces the density of ether groups that are transformed to --OH groups Unglaube et al. (2023). During the reduction process, --CH2-OH groups are also formed beside --OH groups at the diamond surface which approaches the desired surface termination of FNDs Gulka et al. (2024). NV electron spin relaxation $T_{1}$ times of thus modified HPHT FNDs indicate an improvement by $\sim 17$%–$\sim 29$% that were monitored for post-selected single digit FNDs with bright emission and reduced electrical noise Gulka et al. (2024).
These studies inspired us to broaden the characterization of FNDs containing NV centers across a more extensive set of samples towards biological applications.
In our study, we systematically monitor the PL spectra, ODMR spectra, and the $T_{1}$ relaxation times of ensemble NV centers in FNDs as a function of size (10-140 nm) and surface termination. We use commercially available FNDs produced from HPHT diamonds and employ oxygenated and reduced (--OH terminated) samples as references for amino-terminated surfaces. The choice of amino termination (--NH2) facilitates the biological functionalization of FNDs, as --NH2 groups readily enable the attachment of biological molecules Coffinier et al. (2007). However, protonated amino groups can induce a negative electron affinity on diamond surfaces, as observed in CVD diamond surfaces exposed to low-energy NH3 plasma Zhu et al. (2016), which may destabilize NV(--) defects Kawai et al. (2019).
In this work, we employ a novel technique for achieving amino-termination of FNDs based on the Hofmann degradation process. This process consists of multiple wet chemical reaction steps that are milder than plasma treatment, selectively targeting specific oxygen-related chemical groups while also reducing the concentration of sp2 carbon Dai et al. (2016). Notably, we have previously applied this technique to ultrasmall silicon carbide nanoparticles, which share similar oxygen-related surface groups Czene et al. (2023). As a result of the gentle formation of amino-groups on the nanodiamond surface, the modified FNDs may both support stable NV centers and simultaneously serve as functional platforms for biomolecule binding.
Our paper is organized as follows. In Sec. II we describe the materials processing, the experimental tools and methods for characterizing the morphology, size and surface terminations of FNDs as well as the magneto-optical characterization of NV centers in the FNDs. In Sec. III we start the analysis of the FNDs and then we report the results on the PL, ESR and ODMR characterization of the FNDs embedded with NV centers. We discuss the results in the light of previous works here. We conclude our results in Sec. IV. The paper is accompanied by supporting information about the materials processing as well as supporting data as obtained by various experimental methods in the Appendix.
II Methods
FNDs purchased from Adámas Nanotechnologies Inc. were used to study the effects of surface terminations. These FNDs were derived from high-pressure, high-temperature (HPHT) diamonds, which were milled to produce nanodiamonds Shenderova et al. (2019). The nanodiamonds were then electron-irradiated and annealed to generate NV defects. Subsequently, the nanodiamond surfaces were cleaned and oxygenated Shenderova et al. (2019).
Our analysis (see below) revealed that the FND solutions contain sulfur, which may originate from acid cleaning residues or surfactants added to prevent FND aggregation in the solution. Furthermore, post-cleaning of the purchased FND solutions removes some organic sulfur-containing materials (e.g., surfactants or ultrasmall carbon-based particles). However, this process significantly reduces colloidal stability and impacts the examined optical properties, thus we incline to claim that sulfur-related surfactants were present in the as-received FNDs.
The FNDs were either used as is or washed before measurements using solvent exchange centrifugation with DI water three times, acetone two times, and DI water five times, in order to remove the surfactants. We called the initial samples as-received, whereas they were labeled as ”washed” when the surfactants were removed as explained above.
Hydroxyl-terminated FNDs were synthesized via LiAlH4 reduction, adapted from Ref. Shenderova et al., 2011. Briefly, the solvent in the aqueous FND solution was replaced with anhydrous tetrahydrofuran (THF), dried over CaCl2, and removed by centrifugation. Subsequently, 1.6 mL of a 1.0 mol/L LiAlH4 solution was added, resulting in a final LiAlH4 concentration of 0.2 mol/L. The reaction mixture was refluxed overnight under vigorous stirring. Following THF evaporation, the reaction was quenched by adding 1 mol/L HCl. Purification and collection were performed via centrifugation using tubes with polyethersulfone filter membranes. The product was washed multiple times with DI water until a neutral pH was reached, followed by sequential washes with acetone and DI water to ensure complete removal of residual contaminants.
Amino-terminated FNDs were prepared following a modified procedure based on Ref. Czene et al., 2023. Initially, the FND suspension was dried, and SOCl2 was added, followed by sonication at 50 C for 3 hours. The SOCl2 was then removed under vacuum at 50 C, after which an NH3/dioxane solution was introduced. The mixture was sonicated again at 50 C, then heated to 80 C for 10 minutes in the presence of NaOH and Br2. To eliminate excess Br2, NH3 was added to the cooled solution, and any residual NH3 was removed via vacuum evaporation at 50 C. The pH was adjusted using HCl, and purification was carried out through ten cycles of centrifugation and redispersion in water.
We labeled the particles with different types of surface termination as follows: for hydroxyl-termination ”NDx--OH” and for amino-termination,”NDx--NH2” were used where ”x” corresponds to the size of the FND. The considered samples are listed in Table 1.
Table 1: Data of the applied FNDs with giving the sample label, the average diameter of the particle ($d$), the density of the NV centers ($\rho$), the number of NV centers per particle (#NV), and the density of the P1 centers ($\rho_{\text{P}_{1}}$). Data are provided by Adámas Nanotechnologies Co. (Raleigh, NC, USA) and presented in Ref. Shenderova et al., 2019.
Label $d$ (nm) $\rho$ (ppm) #NV $\rho_{\text{P}_{1}}$ (ppm)
 ND140 140 $\leq$ 3 385 $\sim$ 8
ND90 90 $\sim$ 3 200 -
ND70 70 $\leq$ 3 95 $\sim$ 4.8
ND50 50 $\leq$ 3 23 $\sim$ 4
ND40 40 $\leq$ 2 12-14 $\sim$ 2
ND30 30 $\leq$ 2 5-6 -
ND10 10 $<$ 1 1-2 -
Scanning electron microscopy (TESCAN MIRA3 SEM) and high-resolution transmission electron microscopy (Thermo Fisher Themis Cs-corrected HRTEM) were used to observe the shape and size of FNDs. Energy Dispersive Spectroscopy (EDS) spectra were recorded both in SEM and HRTEM to determine the particle composition. For this analysis, the particles were dispersed over a Si substrate in the case of EDS measurements. For the HRTEM analysis, the aqueous suspension of FNDs was drop-cast onto a TEM grid (TED Pella) covered with an ultrathin carbon layer supported by lacey carbon after the purification.
[…]
The photoluminescence (PL) spectra were detected using a Renishaw inVia Raman Microscope with a 50$\times$ Leica objective. The PL spectra were measured on drop-cast FNDs on a Si surface. The excitation wavelength was 532 nm and the power of the laser source was adjusted within 1-800 $\mu$W range and calibrated under the microscope objective. The power range corresponds to 125-2654 $\mu$W/cm2 irradiance. The largest applied power is an upper bound limit in our measurements as a larger power of the laser results in such a high intensity of emission from our FND samples that could be detrimental for the operation of the detectors. For the accuracy of the analysis, we have repeated each measurement 5-7 times on all samples by using the same excitation power. Several measurements were performed using different laser powers on the same area of the samples.
We studied X-band (0.35 T, 9.4 GHz) electron spin resonance (ESR) in a commercial spectrometer (Bruker Elexsys E500). The samples were placed in high quality, defect free quartz tubes. Care was taken to employ a low microwave (MW) exciting power (0.2 mW) and a low magnetic field (0.15-0.25 G) modulation to avoid any distortion to the ESR lineshapes. The measurements were carried out by applying 520-nm laser illumination at room temperature to mimic the conditions of the optically detected magnetic resonance (ODMR) studies. The Matlab-based software EasySpin Stoll and Schweiger (2006) was used to fit the measured ESR data. The P1 center with known spin Hamiltonian parameters was used for calibration of the MW frequency. In the simulation, the $g$ value of the P1 center was fixed, and an anisotropic strain was applied to the hyperfine interaction in order to reproduce the P1 line shape in the spectra. In the initial fitting process, Voigt broadening was tested for the central line; however, Lorentzian broadening was ultimately employed because it provided the best description of the observed peaks. The $g$-factors for the center line were allowed to vary during the fitting procedure.
Continuous wave (cw) ODMR and longitudinal spin relaxation time measurements were conducted at constant excitation power. The zero-field splitting (ZFS) $D$ and $E$ parameters of the observed ensemble NV centers were monitored by cw-ODMR measurements at room temperature (297 K) where $D$ separates the $m_{S}=0$ and the average of the $m_{S}=-1$ and $m_{S}=+1$ levels, $2\times E$ separate the $m_{S}=-1$ and $m_{S}=+1$ levels. In other words, $D-E$ and $D+E$ dips appear in the cw-ODMR spectrum at zero magnetic field. We note that $D=2.87$ GHz and $E=0$ in perfect diamond and non-zero $E$ occurs due to strain or fluctuating electric fields. A 520-nm laser served as the excitation source, delivering 10 mW power before the 0.9 NA Carl Zeiss objective. Emission was collected through the same objective. The collected emission was then passed through a 550-nm dichroic mirror (Semrock) and a 650-nm long-pass filter (Thorlabs) to an APD detector (Excelitas SPCM-AQRH-44). The output of the APD detector was connected to the input of a lock-in amplifier (SR830).
For the microwave field, a Vaunix LabBrick LSG-402 source, a high-power amplifier (Mini-Circuits ZHL-16W-43+), a Mini-Circuits ZASWA-2-50DRA+ switch, and a microwave antenna under the sample were employed. Microwave power ranged from 20 to 40 dBm. The samples were drop-cast onto a non-fluorescent borosilicate cover glass of type No.0 (0.15 mm thickness) and positioned close to the antenna. Evaluation of the results was performed using the EasySpin simulation software Stoll and Schweiger (2006).
In the $T_{1}$ time measurements, the pulse sequence orchestration and data acquisition were performed by Pulse Streamer 8/2 in conjunction with Time Tagger Ultra by Swabian Instruments.
III Results and Discussion
We first report the basic morphological and structural characterization of the FND samples and then we continue the analysis of the surface termination of FNDs upon various treatments. A critical issue is to stabilize the emission from NV(--) because only NV(--) emission from the total PL spectrum can be used for room temperature quantum sensing. For the observation of external electron spins by NV quantum sensor such as radicals in cells leading to diseases Tahirbegi et al. (2016) or to trace chemical reactions evolving radicals Nie et al. (2022), the $T_{1}$ relaxation time should be sufficiently long. To monitor these properties upon surface treatments, we characterize our FNDs by PL, ESR and ODMR techniques where ESR signals may provide insights about the electron spin environment around the NV centers in FNDs.
III.1 SEM analysis
SEM images reveal that the FND particles exhibit an irregular shape with distinct crystal facets in all size ranges (see Appendix B) consistent with prior studies Shenderova et al. (2019). The consequence is that the distance of the NV defects from the nanodiamond surface is most likely much shorter than the half size of the nanodiamond, and the average distance of NV defects from the surface with growing size of nanodiamonds could show a monotonous function but with relatively gentle slope.
III.2 Analysis of the surface termination
We characterized the surface termination of as-received, washed, and modified FNDs by FTIR technique, which was amended by SEM-EDS, HRTEM-EDS and XPS studies on selected FNDs.
Figure 1: Infrared vibration spectra of the as-received, washed, --OH and --NH2 terminated ND10 FNDs.
The FTIR spectra of the as-received samples indicated the presence of oxygen-containing groups that we show on the exemplary ND10 samples in Fig. 1. We note that very similar features were observed for the other FND sizes (Fig. 10 in Appendix C) where we list the detailed analysis in Tables 5-11.
We identified oxygen-related functional groups, including the C--O stretching mode at 1062 $\text{cm}^{-1}$ and a broad band above 3130 $\text{cm}^{-1}$ in the as-received FNDs, originate from O--H stretching in --COOH and --OH groups. The C=O stretching vibrations of carboxyl (--COOH) groups appear at 1773 and 1660 $\text{cm}^{-1}$, providing clear evidence of carboxyl functionalities on the oxygenated diamond surface. These carboxyl groups likely play a dominant role in inducing band bending Freire-Moschovitis et al. (2023). The weak peak at 1332 $\text{cm}^{-1}$ likely corresponds to the C--C stretching vibration of the diamond lattice.
Comparing the FTIR spectra of as-received and washed FNDs reveals two major changes: (i) the C--H bending modes at 1460 $\text{cm}^{-1}$ Petit and Puskar (2018) decrease in intensity after washing, and (ii) the peaks around 1216 $\text{cm}^{-1}$ show a reduction. At the same time, we observed that washed FNDs, regardless of their size, lose their aqueous colloidal stability. This led us to suspect that the vibration bands near 1216 $\text{cm}^{-1}$ correspond to sulfonate groups, which contribute to the stabilization of FND particles in water. However, unambiguous identification from FTIR spectra alone remains challenging due to overlapping C--O stretching vibration bands. To test this hypothesis, we performed SEM-EDS and HRTEM-EDS elemental analyses. The results show that sulfur-related peaks are clearly visible in as-received FNDs but become negligible in the washed samples (see Figs. 8 and 9 in Appendix B). This finding suggests the presence of sulfur-containing organic molecules, or a variety of such molecules, that are weakly attached to the surface. These surfactants play a significant role in colloidal stabilization and, as discussed below, influence the optical properties of FNDs.
The intensity of the C=O stretching vibrations at 1773 and 1660 $\text{cm}^{-1}$ is reduced and the O--H stretching band is narrowed in ND-OH samples (blue curve in Fig. 1), as a result of chemical reduction of the surface groups. It is apparent that the C=O stretching vibration modes did not fully vanish after reduction which implies that the FND surface is not homogeneously terminated with --OH groups. Achieving homogeneity is challenging due to the diverse surface moieties requiring different chemistry for conversion, the unknown effect of facets on the chemical reactions, steric hindrance due to the irregular particle shape, and other factors. Nevertheless, these results are in good agreement with a previous study Shenderova et al. (2011).
For ND-NH2 samples, FTIR spectra showed the reduction of all the oxygen-related peaks and the emergence of nitrogen-related ones (c.f. green and black curves in Fig. 1). Analysis of the FTIR spectra showed the presence of the torsional vibration of the amino-group at 750-800 cm-1 as well as the C-N stretching mode of primary amines, which corresponds to the peaks in the 1060-1090 cm-1 region. The secondary amino-vibration can be found between 1130-1190 cm-1, and the signal of the tertiary amino-groups is between 1160-1210 cm-1. Furthermore, N-H bending vibration band at around 1540-1650 cm-1 also occurs in the FTIR spectra of FNDs after Hofmann degradation (see also Table 5 in Appendix C). This is a good indication that the amino-termination of FNDs was indeed achieved. The same principle applies for amino-terminated FNDs as for the hydroxyl-terminated FNDs, i.e., the surface termination is incomplete, and oxygen-related peaks remain at the surface of FNDs. The presence of primary and secondary amines also indicates inhomogeneity, even for the nitrogen-related groups. XPS studies confirm the effective amino functionalization of FNDs, showing a significant increase in C--N bond intensity relative to C--C bonds after the Hofmann degradation process (Table 2). The recorded XPS spectra and detailed analysis are provided in Appendix D.
Table 2: XPS results on ND10 and ND90 FNDs giving the element concentration of C, N and the ratio of N/C.
Sample C N N/C
ND10 as-received 82.7 1.1 0.013
ND10-NH2 57 1.1 0.019
ND90 as-received 72.5 0 0
ND90-NH2 70.9 0.5 0.007
III.3 Optical properties
Under green illumination, the photoemission spectra of FNDs arises from the superposition of NV(--) and NV(00) emission spectra. Usually, the goal of many FNDs researches is to increase the PL signal of NV(--) which exhibits ODMR signal at room temperature. The typically observed PL spectrum of FNDs is plotted for the as-received ND10 sample in Fig. 2(b). The spectrum can be decomposed to the weighted sum of the PL spectra of NV(--) and NV(00) as plotted in Fig. 2(a). The weights of the two PL signals were determined by the following steps Rondin et al. (2010): (i) the observed PL spectra were normalized to the maximum intensity of the NV(--) phonon sideband, (ii) by taking the known PL spectra of NV(--) and NV(00) in Fig. 2(a), the normalized PL spectra were fitted by the weighted superposition of the PL spectra of N(00) and NV(--) PL as follows
$I(\lambda)=a_{1}\times I_{\text{NV}(-)}(\lambda)+a_{2}\times I_{\text{NV}(0)}(\lambda)\text{,}$ (1)
where $a_{1}$ and $a_{2}$ are the weighting factors. The ratio of fluorescence intensity produced by the NV(--) and the total fluorescence intensity ($f_{\text{NV}}$) calculated as follows
$f_{\text{NV}(-)}=\frac{I_{\text{NV}(-)}}{I_{\text{NV}(-)}+I_{\text{NV}(0)}}\text{.}$ (2)
Figure 2: Photoluminescence spectrum analysis of FNDs. (a) Normalized photoluminescence spectra of NV(00) and NV(--) defects. (b) Photoluminescence spectrum of the as-received ND10 sample, where the fitted curve is the weighted superposition of NV(00) and NV(--) photoluminescence spectra.
Figure 3: Photoluminescence spectrum analysis of FNDs. NV(--) emission fraction for the (a) washed, (b) --OH terminated, and (c) --NH2 terminated FNDs as a function of the applied laser intensity for different FND sizes.
There are factors, intrinsic and extrinsic to the NV defect in diamond that determine the observed $f_{\text{NV}(-)}$ values. Strong illumination of NV defect with a green laser will continuously cycle between NV(--) and NV(00) (see Ref. Gaebel et al., 2006) because of two-phonon ionization with successive steps via the optical excited state (see Ref. Gali, 2019 and references therein). We note that understanding the microscopic mechanism of this process is still under intense research Razinkovas et al. (2021); Wirtitsch et al. (2023); Thiering and Gali (2024). This is a non-linear optical process that depends on the optical excitation power. Illumination may also ionize other defects in diamonds, such as the nitrogen donor, by ejecting an electron and vacancy clusters with ejecting holes, which depends on the doping concentration of nitrogen and the formation of NV defects via irradiation and annealing. We note that the latter process occurs for neutral vacancy clusters, which form due to band bending in near-surface regions Neethirajan et al. (2023); Freire-Moschovitis et al. (2023). Surface defects are typically acceptor-like Kaviani et al. (2014); Stacey et al. (2019); Dwyer et al. (2022); Chou et al. (2023); Dontschuk et al. (2023) that also eject holes after photo-ionization. As a consequence, the NV(--) emission from the total emission is a photodynamics process, which depends on the excitation power, defect concentrations inside the FNDs, band bending from the surface towards the core of FNDs, and density of the surface defects on FNDs.
We plot the observed $f_{\text{NV}(-)}$ values as a function of size, surface termination, and laser intensity in Fig. 3 where we show the data with the least and most intense laser powers (see other results in Appendix E). The $f_{\text{NV}(-)}$ of the as-received samples increased with particle size, following a power law and reaching a plateau at approximately 70 nm in diameter. This trend aligns with previous results reported for single NV(--) centers Rondin et al. (2010) and ensembles Wilson et al. (2019). For all other studied FNDs, the trend shifted, with the plateau occurring at smaller sizes: 40 nm for ND-OH and 20 nm for both washed and ND-NH2 samples. This phenomenon may be understood by assuming that surface acceptor states can be photo-ionized with ejecting holes that would convert NV(--) to NV(00). The scattering $f_{\text{NV}(-)}$ data can be understood as the distances of NV defects from the diamond surface are not necessarily longer with larger FND size because of the irregular shape of FNDs. Another general trend is that increasing the power of the laser leads to larger $f_{\text{NV}(-)}$ values for the smallest FNDs (ND10) except for washed FNDs to which $f_{\text{NV}(-)}$ remains at $\sim 0.3$. The origin of this phenomenon is unclear as $f_{\text{NV}(-)}$ goes above 0.4 for as-received small FNDs (see Fig. 12 in Appendix E); nevertheless, one may argue that intense laser irradiation may induce structural changes at the surface that remove acceptor states that has the most significant effect for the smallest FNDs. We do not have such an evidence from FTIR studies that might be understood with the fact these surface acceptor defects could be topographically protected and chemically stable at the interface of facets Chou et al. (2023). For the medium sized FNDs (50-90 nm) the $f_{\text{NV}(-)}$ values typically scatter between 0.5 and 0.6. The largest $f_{\text{NV}(-)}$ values occur for --NH2 terminated ND140 FNDs at around 0.8 while it stays around 0.6 for as-received, washed or --OH terminated FNDs at the same size. This clearly demonstrates that the surface termination affects the properties of NV centers embedded in relatively large FNDs. FTIR analysis does not reveal any obvious reason for the outstanding $f_{\text{NV}(-)}$ value for --NH2 terminated ND140 FNDs as the FTIR spectra of --NH2 FNDs exhibit similar features and characteristics in various sizes apart from small shifts (see Tables 5-11 in Appendix C).
III.4 Paramagnetic centers in FNDs
The photostability of NV(--) in FNDs depends on the nearby acceptor and donor defects that could be paramagnetic; these paramagnetic species can be monitored by ESR spectroscopy. We performed electron spin resonance measurements on selected sample sizes (ND30, ND70, ND90, ND140), to characterize the paramagnetic centers in FNDs. The paramagnetic centers also strongly influence the $T_{1}$ time of embedded NV centers as explained in the Introduction. The $T_{1}$ can be observed under green illumination which changes the charge state of defects affecting their spin states, thus we simulated this environment by photo-ESR studies where the FND samples were irradiated by 520-nm laser during ESR measurements.
Since the free amino-group is unstable in air, the measurements were performed in a liquid medium, 1,4-dioxane having a low dielectric constant of 2.25. The ESR spectra for the studied FNDs are plotted in Fig. 4.
Figure 4: ESR spectra and fitted plots for FNDs. Washed and --NH2 terminated FNDs are shown in the left and righ columns, respectively.
In agreement with previous findings Shenderova et al. (2019); Shames et al. (2015); Panich et al. (2015), the ESR spectra exhibit a central line that is the superposition of two Lorentzian signals: a broader signal attributed to carbon-inherited paramagnetic centers (primarily unpaired $S=1/2$ electron spins of dangling bonds formed during diamond milling), the P1 center that appears as a 14N hyperfine pattern broadened by the inhomogeneous strain field of nanodiamonds and a narrow $S=1/2$ ESR center. The origin of the narrow $S=1/2$ ESR center is not known. We hypothesize that it could be the ESR signal of paramagnetic vacancy clusters inside the nanodiamonds or close to the surface of nanodiamonds that are created by electron irradiation and annealing to mediate the conversion of substitutional nitrogen defects to NV defects.
The main spin Hamiltonian parameters are listed in Table 3. The 14N hyperfine pattern creates a visible signal outside of the broad central line at around 333 and 339 mT where the central line is dominated by the surface-related $S=1/2$ species in ND90 and ND140 samples. The signal of P1 center is masked by the other $S=1/2$ species for nanodiamonds smaller than 90 nm, although, it is weakly observable for ND70. Nevertheless, it can be concluded from the analysis of the ESR spectra of ND90 and ND140 samples that the signal of P1 center has higher relative intensity in amino-terminated FNDs than that in as-received FNDs. Since surface modification reactions affect only the outer moieties (thus altering the surface-related dangling bonds’ concentration without impacting P1 centers), the observed increase in the relative intensity of the P1 center suggests a reduction of surface-related defects due to the Hofmann degradation process. Interestingly, the narrow $S=1/2$ ESR center does not show a clear trend upon Hofmann degradation of ND90 and ND140 samples. For ND90 samples, the relative concentration of the narrow $S=1/2$ ESR center decreases after Hofmann degradation while it increases for ND140 samples. This result implies that if the narrow $S=1/2$ ESR defects reside close to the surface of nanodiamonds they are not directly affected by the surface modifications. This might be a complex interplay of various effects: as the number of dangling bonds decreases at the surface upon Hofmann degradation process the band bending may incline downwards due to --NH2 groups but the near-surface substitutional nitrogen donors could easily transfer their electrons towards the acceptor-like near surface vacancies Peng et al. (2019) that becomes paramagnetic. With assuming the charge stabilization of the paramagnetic defects near the NV center in ND140 samples after Hofmann degradation process, the outstanding $f_{\text{NV}(-)}$ might be explained that does not occur for smaller FNDs.
Table 3: Simulated electron spin resonance spectra with using three paramagnetic defects. The main spin Hamiltonian parameters are listed where the $g$ factor of the P1 center is fixed at the experimental data Smith et al. (1959); Shames et al. (2015); Panich et al. (2015). The ESR parameters determined for the P1 centers by simulation are $g_{\mathrm{iso}}=2.0024$, $A_{zz}=115$ MHz and $A_{xx}=A_{yy}=82$ MHz. The individual line widths of $\Delta H_{\mathrm{pp}}^{\mathrm{L}}$ with Lorentzian function and the strain applied in the fitting procedure of the hyperfine interaction are shown in MHz unit. The $A_{\mathrm{strain}}$ values are for $A_{zz}$ and $A_{xx}=A_{yy}$, respectively.
P1 center ($S=1/2$) Broad center ($S=1/2$) Narrow center ($S=1/2$)
Sample $g$ $\Delta H_{\mathrm{pp}}^{\mathrm{L}}$ $A_{\mathrm{strain}}$ Weight $g$ $\Delta H_{\mathrm{pp}}^{\mathrm{L}}$ Weight $g$ $\Delta H_{\mathrm{pp}}^{\mathrm{L}}$ Weight
ND30 - - - 0.0000 2.0030(2) 1.002 0.8749 2.0029(2) 0.006 0.1251
ND30-NH2 - - - 0.0000 2.0030(2) 1.187 0.9231 2.0029(2) 0.349 0.0768
ND70 2.00240 0.710 10.68, 7.03 0.0022(18) 2.0030(2) 1.174 0.9103 2.0028(2) 0.388 0.0875
ND70-NH2 2.00240 0.775 10.68, 5.54 0.0024(16) 2.0030(2) 1.61 0.9109 2.0028(2) 0.712 0.0867
ND90 2.00240 0.691 10.68, 7.35 0.0257 2.0028(2) 1.103 0.7665 2.0027(2) 0.306 0.2078
ND90-NH2 2.00240 0.717 10.68, 5.21 0.0768 2.0030(2) 1.261 0.7473 2.0029(2) 0.348 0.1760
ND140 2.00240 0.419 10.68, 3.01 0.1074 2.0030(2) 1.308 0.7503 2.0029(2) 0.356 0.1423
ND140-NH2 2.00240 0.481 10.68, 1.65 0.1185 2.0030(2) 1.457 0.7265 2.0029(2) 0.362 0.1550
III.5 Optically detected magnetic resonance spectrum and longitudonal spin relaxation time
The quantum sensor operation of FNDs relies on the ODMR signals of NV(--) defects. This technique combines photoluminescence with electron spin resonance. In our study, we record the ODMR signals for the ensemble of FNDs. The intensity of fluorescence was relatively weak for the smallest FNDs in our study (30 nm). For those samples, we picked up the brightest spots that we found in the confocal map of the samples. A typical cw-ODMR signal is plotted for the representative ND50-OH FNDs in Fig 5(a). The thin black curve is the recorded ODMR data whereas the red curve is the simulation of the ODMR signal with the use of spin-Hamiltonian of NV(--) defect with allowing inhomogeneous broadening. Apparently, a double negative peak arises which is a signature of breaking the $C_{3v}$ symmetry of the defect. The small valley between the two peaks is the $D$ parameter whereas the two peaks are separated by $2\times E$. We recorded the ODMR signal for every type of FNDs considered in our study and fit the results to extract the $D$ and $E$ parameters that we list in Table 4. We find a general trend that the $E$ parameters shifts from $\sim 8$ MHz towards $\sim 5$ MHz as we increase the size of FNDs from 30 nm to 140 nm, independently from the surface termination. The size dependence of the observed $D$ parameters differs: it shows up $\sim 2867$ MHz in 10-nm sized FNDs and then it increases to $\sim 2870$ MHz for larger FNDs (40 nm) but then the observed $D$ parameters fluctuate around this value for larger FNDs.
Figure 5: ODMR results on ND50-OH. (a) cw-ODMR and (b) $T_{1}$-relaxometry measurements of ND50-OH sample. The microwave $\pi$ pulse is typically applied for 500 ns. Red curves are simulated data.
There could be two reasons for the symmetry breaking $E$ splitting in the ODMR spectrum: static strain and (fluctuating) electric fields. Theoretical analysis and simulations showed that two fields dissimilarly affect the ODMR spectrum on ensemble of NV defects: static strain both shifts the $D$ parameter and create an $E$ splitting whereas the electric fields do not change the $D$ parameter but only create $E$ splitting Mittiga et al. (2018). The data in Table 4 imply that the static strain affects the spin levels of the electronic ground state in 10-nm and 30-nm sized FNDs whereas it has minor effect in larger FNDs as the $D$ parameters do not shift in the latter. According to a recent ab initio study, the $E$ parameter fast converges below 5 MHz with increasing the distance between NV(--) and the (100) facet of the diamond surface Pershin et al. (2025). For larger nanodiamonds, it is more likely that NV centers reside farther to the diamond surface than 2 nm, thus our conclusion is well supported by this study. The size dependent change in the $E$ parameter of the spin levels in the NV centers’ ground state embedded in larger ($>30$ nm sized) FNDs should be associated with the electric field noise. The electric field noise around the NV centers is typically caused by charging the parasitic defects of diamond with illumination that is used to optically spin-polarize or read out the spin state. The source of the parasitic defects could be the substitutional nitrogen donor defects, acceptor-like vacancy-cluster defects around the NV center and acceptor-like near-surface or surface defects. The distance between the surface defects and the NV center increases in larger FNDs that reduces the value of the $E$ parameter. This explanation is consistent with the interpretation of the ESR data where the density of the broad $S=1/2$ ESR centers associated with surface dangling bonds decreases either with increasing size of FNDs or going from the as-received to amino-terminated samples with a given size of FNDs, and the respective $E$ parameters follow the same trend.
Table 4: Magnetic resonance parameters for various investigated FNDs.
NDx as-received --OH --NH2
$d$ (nm) $D$ (MHz) $E$ (MHz) $T_{1}$ ($\mu$s) $D$ (MHz) $E$ (MHz) $T_{1}$ ($\mu$s) $D$ (MHz) $E$ (MHz) $T_{1}$ ($\mu$s)
10 2867.5 8.0 (8 $\pm$ 1.2) 2865.7 8.6 (11 $\pm$ 4.1) 2867.5 8.0 (31 $\pm$ 7.0)
30 2868.5 7.3 (25 $\pm$ 3.6) 2869.0 7.2 (39 $\pm$ 6.6) 2868.5 7.1 (21 $\pm$ 2.8)
40 2869.4 7.3 (32 $\pm$ 1.7) 2868.7 5.8 (34 $\pm$ 12.0) 2870.2 6.8 (16 $\pm$ 1.0)
50 2870.0 7.3 (27 $\pm$ 4.9) 2870.0 5.5 (46 $\pm$ 7.2) 2869.7 6.8 (18 $\pm$ 1.1)
70 2870.0 6.8 (32 $\pm$ 3.0) 2870.5 6.2 (40 $\pm$ 4.0) 2870.5 6.0 (42 $\pm$ 3.7)
90 2870.5 6.4 (64 $\pm$ 5.6) 2869.5 5.6 (38 $\pm$ 3.7) 2869.4 5.7 (24 $\pm$ 2.8)
140 2870.3 5.7 (74 $\pm$ 6.7) 2870.5 4.5 (83 $\pm$ 9.7) 2870.0 5.0 (32 $\pm$ 2.8)
As a next step we observed the longitudonal spin relaxation time of ensemble of NV centers in FNDs with the protocol as the inset depicts in Fig. 5(b). The NV centers were initially polarized into the $m_{s}=0$ state by an optical pulse lasting 4 $\mu$s. Next, a microwave $\pi$ pulse was applied, typically for 500 ns. Next, after the free evolution time $\tau$, a second optical pulse is applied for the readout ($R$) lasting 3 $\mu$s. Then the same sequence is applied but with the $\pi$ pulse omitted, and the resulting time $T_{1}$ is derived from the $1-R(\pi_{\text{on}})/R(\pi_{\text{off}})$ function exponential decay.
A typical experimental curve is plotted for ND50-OH sample (thin black line) whereas the red curve is a single exponential decay function fit (red curve) to the experimental data in Fig. 5(b). We did not apply any post-selection on the FNDs but we simply conducted these measurements on those samples that were sufficiently bright and exhibited well-detectable cw-ODMR spectrum. The observed $T_{1}$ times are listed for all the considered as-received, --OH terminated and --NH2 terminated FNDs in Table 4. The observed logitudonal spin relaxation rate, $1/T_{1}$ for a few selected washed FNDs are plotted in Fig. 6, together with the all the other observed $1/T_{1}$ values. The $1/T_{1}$ rates for the smallest ND10 samples are high and the uncertainties in the exponential fit is relatively large because the $T_{1}$ times are comparable to the duration of the optical pulses used to spin-polarize and read out the electron spin of NV centers. Notably, the amino-terminated ND10, ND10-NH2, is an exception for which the observed $T_{1}$ time is significantly longer than that of the other ND10 samples. This might be attributed to the chemical treatments in the Hofmann degradation process that could be effectively remove $sp^{2}$ type of dangling bonds Dai et al. (2016) unlike the other methods. Unfortunately, the dangling bonds are not directly visible in the FTIR spectra and the ESR signal from ND10 samples was too weak for quantitative analysis, thus this assumption cannot be supported by experimental data in our hand. In general, the trends on the $1/T_{1}$ rates are distinct for the --NH2 terminated FNDs and the other FNDs. For the oxygenated and --OH terminated FNDs the $1/T_{1}$ rate slows down with growing size of FNDs, although this is not a smooth monotonous function. On the other hand, the $1/T_{1}$ rate scatters around $4\times 10^{4}$ Hz, i.e., the $T_{1}\approx 25$ $\mu$s, for all amino-terminated FNDs in the considered size range. Nevertheless, the advantage of --NH2 termination of the smallest FNDs is not transferable to larger ones, e.g., the oxygenated and --OH terminated ND140 samples have about by half slower $1/T_{1}$ rates than that of amino-terminated ND140.
In our samples, the concentration of P1 centers is relatively high and contains 13C nuclear spins with natural abundance, therefore, it is expected that the $T_{1}$ times of the NV center will be much shorter than 1 ms. The $T_{1}$ times can be reduced due to electron spin flip-flop processes that were well observable in special bulk diamond samples with dense NV defects Choi et al. (2017), and it was analyzed by ab initio calculations and found that acceptor defects with a distance of about 4 nm results in spin flip-flop processes with a rate comparable with the $T_{1}$ relaxation times of NV center at room temperature Chou et al. (2018). As the surface states and near-surface states are acceptor defects a similar electron spin flip-flop process can occur between the embedded NV center and these defects in FNDs. As a consequence, a general trend is expected that the $T_{1}$ time of NV centers will be longer with increasing size of FNDs because that significantly reduces the electron spin flip-flop rates. This trend is indeed observable for oxygenated and --OH terminated samples. We note again that the shape of our FNDs is not spherical, thus the non-monotonous change in the $T_{1}$ times with increasing size of FNDs may be explained so that the average distance between the NV center and the FND surface does not proportionally change with size of our FNDs. On the other hand, the amino-terminated FNDs do not show this global trend anticipated above. By combining the PL and ESR studies on amino-terminated FNDs, we concluded that the negative charge state of near-surface defects could be stabilized by --NH2 groups but those negatively charged near-surface defects are paramagnetic. The --NH2 terminated FNDs exhibit more negative electron affinity than --OH terminated or oxygenated FNDs. This could affect the charge state of the deeply buried vacancy-like defect in FNDs that are near NV centers. If those charge states of defects stabilize paramagnetic ground state then the electron spin flip-flop process will be effective and could shorten the $T_{1}$ time of NV centers. Nevertheless, the observed $T_{1}$ times in amino-terminated FNDs fall the same order magnitude as those of oxygenated and --OH terminated FNDs.
Figure 6: Longitudinal spin-relaxation rate (1/$T_{1}$) of the FNDs as a function of the FND size. The line is a guide for the eyes.
IV Conclusion
We studied the fundamental properties of FNDs embedding NV centers, including particle shape, the vibrational spectrum of surface groups, the PL spectrum of NV defects, and paramagnetic defects under illumination, which are also used to detect the ODMR signal from NV centers. Additionally, we examined longitudinal spin relaxation times ($T_{1}$) as a function of FND surface termination.
In particular, we applied the Hofmann degradation process to achieve amino termination of FNDs, a modification beneficial for biological applications. Among these, one of the most prominent uses of FNDs is electron spin relaxometry, which can be employed, for instance, to monitor radical formation in living cells due to inflammation. Our results demonstrate that the Hofmann degradation process effectively modifies the surface termination by removing surface defect spins and partially introducing --NH2 groups directly linked to the diamond surface.
Interestingly, while --NH2 groups shorten $T_{1}$ relaxation times for larger FNDs, they extend $T_{1}$ relaxation times for smaller FNDs (e.g., 10 nm-sized) compared to oxygenated or --OH terminated FNDs. This behavior is attributed to the removal of sp2 dangling bonds at the FND surface during Hofmann degradation, while maintaining a lower electron affinity compared to oxygenated and --OH terminated FNDs. Nevertheless, amination via Hofmann degradation leaves certain oxygen-related groups intact on the FND surface, allowing for stable NV centers.
We conclude that the prepared amino-terminated FNDs can be simultaneously functionalized with desired biological molecules and employed for electron spin relaxometry measurements, making them an attractive quantum sensor probe.
Acknowledgements.
The work was supported by the Quantum Information National Laboratory sponsored by National Research, Development and Innovation Fund (NKFIH) grant no. 2022-2.1.1-NL-2022-00004 as well as by grant no. VEKOP-2.3.3-15-2016-00002 of the European Structural and Investment Funds. The research reported in this paper and carried out at Wigner Research Centre for Physics is supported by the infrastructure of the Hungarian Academy of Sciences.
Author Contributions
Conceptualization, N.J., D.B. and A.G.; methodology, N.J., O.K., Zs.Cz., D.B. and A.G.; investigation, N.J., V.V., O.K., Zs.Cz., Sz.Cz. and A.Cs.; project administration, A.G.; resources, A.G.; supervision, A.G.; writing—original draft, N.J. and A.G.; writing—review and editing, N.J. and A.G. All authors have read and agreed to the published version of the manuscript.
Appendix A Sample preparation
In order to obtain the pure diamond particles the NPs were washed six times by the cycle of centrifugation and suspension in water. As the cleaning procedure finished the particles became even more prone to aggregation, so ultrasonication had to be applied before each measurement.
To realize the --OH termination the following procedure was used. The FND particles were aggregated in Eppendorf centrifuge tubes then removing the supernatant Tetrahydrofuran (THF) - dried with CaCl2 - was added and the mixture was dispersed. This step removed the surfactant from the FNDs. In a round bottom flask the redispersed FND solution was diluted with 5 mL CaCl2 dried THF and added 1.6 mL of 1.0 mol LiAlH4 in THF solution. This reaction mixture was refluxed intensely overnight. After evaporating THF, the reaction mixture was quenched by adding 1 mol HCl. The cleaning and collecting took place by centrifugation using centrifuge tubes with polyethersulfone filter media. The product was washed with Millipore water several times to achieve a neutral pH condition. Finally, it was washed with acetone several times.
The amino-terminated FND particles were created by applying Hofmann-degradation. First, the FND suspension was dried. SOCl2 was added to the FNDs and sonicated at 50 C for 3 hours. The SOCl2 was removed under vacuum at 50 C and NH3/dioxane solution was added to the FNDs. This was followed by a new sequence of sonication at 50 C. The solution with NaOH and Br2 was heated at 80 C for 10 minutes. NH3 is added to the cooled mixture to remove the Br2 excess. Finally, NH3 was eliminated by evaporation in vacuum at 50 C. The pH value was set by applying HCl. The solution was cleaned by repeating the cycle of the centrifugation and suspension in water 10 times. The surfactant (since the original suspension of FND contained sulfur) was removed at the final purification step.
Appendix B Analysis by SEM, HRTEM, and EDS
The SEM images clearly show the non-spherical shape of the FNDs in Fig. 7. Fig. 8 presents the EDS spectra of ND10 as-received and washed samples as obtained by SEM. The presence of sulfur is evident in the as-received samples. This observation was also proven by the EDS analysis performed in HRTEM (Fig. 9).
Figure 7: SEM image of ND90 ensemble and individual particles. The white line corresponds to 200 nm.
Figure 8: EDS spectra of (a) ND10 as-received and (b) ND10 washed particles.
Figure 9: HRTEM image and HRTEM-EDS analysis of (a) ND10 as-received and (b) ND10 washed particles. The scale corresponds to 10 nm. The black line indicates the measurement of the size of the nanodiamond particles.
Appendix C FTIR supporting data
In Tables 5-11, a detailed list of vibration bands of the FTIR spectra (presented in the respective Figs. 1 and Figs. 10) where the association of the observed vibration bands with the respective chemical groups at the surface are summarized. The FTIR analysis was done on as-received, washed, hydroxyl, and amino-terminated samples for ND10, ND50, and ND90 FNDs, while for ND30, ND40, ND70, and ND140 FNDs we did not study the washed samples. The FTIR spectra of ND40 were really similar to ND30, therefore it is not shown in detail.
Table 5: FTIR data measured on the ND10 FNDs with various surface terminations. The observed vibration bands, the associated functional groups, and the respective vibration modes are given.
band (cm-1) functional groups vibration modes
ND10 as-received
3590-3130 --COOH, --OH O-H stretching
2960-2856 --CH3, --CH2 C-H stretching
1850-1660 --COOH, --CO, --CHO, C=C C=O from carboxyl group (monomer, dimer), acid-anhydride (symmetric and asymmetric), H bond in chelate bond, C=C stretching
1440 --CH3, C-C bending mode, perpendicular to the plane, C-C skeletal
1270-1180 --SO3X, C-O-C from sulfonate, symmetric and asymmetric C-O-C stretching
1180-1000 --OH C-O stretching modes (primary, secondary, tertiary)
ND10 --OH terminated
3600-3120 --OH O-H stretching
2953-2850 --CH3, --CH2 C-H stretching
1660 --C=O C=O from unreacted COOH group, C=C stretching
1463 --CH3 bending mode, perpendicular to the plane
1240-1000 --C-O, C-C C-O stretching modes (primary, secondary, tertiary), skeletal vibration
ND10 --NH2 terminated
3320 --NH2, --NH N-H stretching
2960-2855 --CH3, --CH2 C-H stretching
1740 --CHO, --C=O C=O stretching
1670 --CHO ortho-hydroxy, amino and formyl group, chelate H bond
1650-1540 --NH2, NH N-H bending
1467 --CH3 bending mode, perpendicular to the plane
1390 --C-N aniline-like and tertiary C-N stretching
1190-1050 --C-N secondary, primary C-N stretching
Table 6: FTIR data measured on the ND30 FNDs with various surface terminations. The observed vibration bands, the associated functional groups, and the respective vibration modes are given.
band (cm-1) functional groups vibration modes
ND30 as-received
3550-3090 --COOH, --OH -OH stretching
2980-2815 --CH3, --CH2 C-H stretching
1860-1560 --CO, --CHO, --OH C=O from carboxyl group (monomer, dimer), acid-anhydride (symmetric and asymmetric), -OH in intramolecular H-bond, chelate bond
1471 --CH3, C-C bending mode, perpendicular to the plane, C-C skeletal
1300-1000 --SO3X, C-O-C, --C-O from sulfonate, symmetric and asymmetric C-O-C stretching C-O stretching (primary, secondary, tertiary)
ND30 --OH terminated
3500-3120 --OH O-H stretching
2980-2790 --CH3, --CH2 C-H stretching
2100-1930 --OH in intramolecular H bond, chelate bond
1820-1700 --CHO, --CO unreacted -CHO and -CO groups, chelate H bond
1680-1550 --CHO, --CO C=O stretching, unreacted -CHO and -OH in chelate H bond (in ortho position next to each other), adsorbed water
1450 --CH3 bending mode, perpendicular to the plane
1300-1100 --C-O, C-C C-O stretching modes (primary, secondary, tertiary), skeletal vibration
710 --OH O-H deformation mode in H-bond
ND30 --NH2 terminated
3550-3300 --NH2, --NH N-H stretching
2980-2825 --CH3, --CH2 C-H stretching
1733 --NH2CO, --NHCO unreacted amide groups ( cyclic amide-lactam)
1640-1560 --NH2, --NH N-H bending
1450 --CH3 bending mode, perpendicular to the plane
1390-1050 --C-N aniline-like and tertiary, secondary, primary C-N stretching
705 --NH2 N-H bending perpendicular to the plane
Table 7: FTIR data measured on the ND40 FNDs with various surface terminations. The observed vibration bands, the associated functional groups, and the respective vibration modes are given.
band (cm-1) functional groups vibration modes
ND40 as-received
3600-3010 --COOH, --OH O-H stretching
2970-2840 --CH3, --CH2 C-H stretching
1860-1620 --CO, --CHO, --OH C=O from carboxyl group (monomer, dimer), acid-anhydride (symmetric and asymmetric), OH in intramolecular H-bond, chelate bond
1473 --CH3, C-C bending mode, perpendicular to the plane, C-C skeletal
1350 --SO3X, C-O-C, --C-O from sulfonate, symmetric and asymmetric C-O-C stretching C-O stretching (primary, secondary, tertiary)
ND40 --OH terminated
3520-3100 --OH O-H stretching
2980-2790 --CH3, --CH2 C-H stretching
1720-1600 --CHO, --CO unreacted -CHO and -CO groups, chelate H bond
1680-1550 --CHO, --CO, C=C, --OH C=O stretching,-unreacted -CHO and -OH in chelate H bond (in ortho position next to each other), adsorbed water, C=C stretching
1450 --CH3 bending mode, perpendicular to the plane
1360 --OH O-H bending
1290-1220 C-O-C, --C-O C-O stretching (secondary, tertiary)
1080-1000 --C-O C-O stretching (primary, secondary)
720 --OH O-H deformation mode in H-bond
ND40 --NH2 terminated
3540-3090 --NH2, --NH N-H stretching
2980-2830 --CH3, --CH2 C-H stretching
1845-1580 --NH2CO, --NHCO, --NH2, --NH unreacted amide groups ( cyclic amide-lactam), N-H bending
1450 --CH3 bending mode, perpendicular to the plane
1410 amide III band C-N stretching, NH2 bending, C=O stretching
1320-1000 --C-N aniline-like and tertiary, secondary, primary C-N stretching
806 --NH2 N-H bending perpendicular to the plane
Table 8: FTIR data measured on the ND50 FNDs with various surface terminations. The observed vibration bands, the associated functional groups, and the respective vibration modes are given.
band (cm-1) functional groups vibration modes
ND50 as-received
3450-3000 --COOH, --OH O-H stretching
2925, 2855 --CH3, --CH2 C-H stretching
1835-1605 --COOH, --CO, --CHO, C=C C=O from carboxyl group (monomer, dimer), acid-anhydride (symmetric and asymmetric), H bond in chelate bond, C=C stretching
1522 --C=O, C=C, C-C skeletal vibrations, -C=O from diketone
1270-1130 --C-O tertiary, secondary and aryl like C-O, asym.-symm. C-O-C
1390 --C-N aniline-like and tertiary C-N stretching
1190-1050 --C-N secondary, primary C-N stretching
890-835 C-O-C asymmetric,symmetric stretching from dioxane and epoxid like C-O-C units
ND50 --OH terminated
3600-3000 --OH O-H stretching
2928, 2855 --CH3, --CH2 C-H stretching
1230-1000 --C-O tertiary, secondary and aryl like C-O
ND50 --NH2 terminated
3250 --NH2 N-H associated form
1760-1680 --CHO, --C=O ortho-hydroxy, amino and formyl group, chelate H bond
1220-1080 C-N C-N stretching
800 --NH2 N-H bending mode, perpendicular to the plane
Table 9: FTIR data measured on the ND70 FNDs with various surface terminations. The observed vibration bands, the associated functional groups, and the respective vibration modes are given.
band (cm-1) functional groups vibration modes
ND70 as-received
3530-3090 --COOH, --OH O-H stretching
2980-2830 --CH3, --CH2 C-H stretching
1980-1800 --OH, --CO, --CHO C=O from carboxyl group (monomer, dimer), acid-anhydride (symmetric and asymmetric), -OH in intramolecular H-bond
1800-1600 --CO, --CHO, --OH, C=C C=O from carboxyl group (monomer, dimer), acid-anhydride (symmetric and asymmetric), -OH in intramolecular H-bond, chelate bond, C=C stretching
1455 --CH3 bending mode, perpendicular to the plane, C-C skeletal
1220-1010 --SO3X, C-O-C, --C-O from sulfonate, symmetric and asymmetric C-O-C stretching C-O stretching (primary, secondary, tertiary)
ND70 --OH terminated
3580-3110 --OH O-H stretching
2990-2830 --CH3, --CH2 C-H stretching
1980-1800 --OH, --CO, --CHO C=O from carboxyl group (monomer, dimer), acid-anhydride (symmetric and asymmetric), -OH in intramolecular H-bond
1730-1550 --CH3, --CH2 C=O stretching,-unreacted -CHO, CO and -OH in chelate H bond (in ortho position next to each other), adsorbed water, C=C stretching
1480, 1440 --OH, --CO, --CHO bending mode, perpendicular to the plane
1300 --OH O-H bending
1250, 1160, 1100 --C-O C-O stretching (primary, secondary, tertiary)
710 --OH O-H deformation mode in H-bond
ND70 --NH2 terminated
3540-3070 --NH2, --NH N-H stretching
2975-2800 --CH3, --CH2 ortho-hydroxy, amino and formyl group, chelate H bond
1770-1490 --NH2CO, --NHCO, --NH2, --NH unreacted amide groups ( cyclic amide-lactam), N-H bending
1460-1300 --CH3, --CH2, C-N C-H bending mode, perpendicular to the plane, C-N stretching (aniline-like)
1250 --C-N tertiary, secondary C-N stretching
1080-1020 --C-N secondary, primary C-N stretching
795 --NH2 N-H bending perpendicular to the plane
Table 10: FTIR data measured on the ND90 FNDs with various surface terminations. The observed vibration bands, the associated functional groups, and the respective vibration modes are given.
band (cm-1) functional groups vibration modes
ND90 as-received
3380-3050 --COOH, --OH O-H stretching
1790 --C=O C=O from acid anhydride, symmetric and antisymmetric
1690-1600 --C=O C=O from carboxyl group
1480 C-C C-C skeletal vibrations
1310 --C-O, S-O symm. C-O stretching (-CO-O) and O-H bending from sulfonate
1110 --C-O, C-O-C C-O stretching from C-OH, anhydride and ether-like
ND90 --OH terminated
3600-3070 --OH O-H stretching
2960-2845 --CH3, --CH2 it can be the CO of an unreacted COOH group, or e.g., chelate bond seen by the ortho OH-COOH group (ortho-next to each other), $\beta$-diketo carbonyl moiety
1110 --C-O C-O stretching mostly from secondary alcoholic group
ND90 --NH2 terminated
3370 NH2 N-H stretching
2970-2855 --CH3, --CH2 C-H stretching
1785 --CHO/--C=O ortho-hydroxy/amino and formyl group, chelate H bond
1650-1560 --NH2, NH primary and secondary N-H bending mode
1460 --CH3, C-C CH3 bending mode perpendicular to the plane, C-C skeletal mode
1320 C-N aromatic-like and secondary C-N stretching
1212-1100 C-N primary and secondary C-N stretching
808 NH2 N-H bending mode, perpendicular to the plane
Table 11: FTIR data measured on the ND140 FNDs with various surface terminations. The observed vibration bands, the associated functional groups, and the respective vibration modes are given.
band (cm-1) functional groups vibration modes
ND140 as-received
3600-3060 ---COOH/--OH O-H stretching
2980-2830 --CH3/--CH2 C-H stretching
1820-1620 --COOH/--CO/--CHO C=O from carboxyl group (monomer, dimer), acid-anhydride (symmetric and asymmetric)
1460 --CH3, C-C bending mode, perpendicular to the plane, C-C skeletal
1350-1120 --SO3X, C-O-C, --C-O from sulfonate, symmetric and asymmetric C-O-C stretching C-O stretching (secondary, tertiary)
1180-1000 C-O C-O stretching modes (primary, secondary, tertiary)
ND140 --OH terminated
3550-3100 --OH O-H stretching
2980-2830 --CH3/--CH2 C-H stretching
1740-1600 --COOH/--CO/--CHO C=O from unreacted carboxyl group (monomer, dimer), -CHO formyl group from reduced -COOH (C=O symmetric and asymmetric stretching))
1460 --CH3 bending mode, perpendicular to the plane
1150-980 --C-O, C-C C-O stretching modes (primary, secondary, tertiary), skeletal vibration
ND140 --NH2 terminated
3500-3080 -NH2/--NH N-H stretching
2960-2790 --CH3/--CH2 C-H stretching
1730 --C=O C=O stretching from unreacted amide (lactam, cyclic amide)
1670 --CHO, NH ortho-hydroxy/amino and formyl group, chelate H bond
1580-1540 --NH2/NH N-H bending
1457 --CH3 bending mode, perpendicular to the plane
1390-1160 --C-N aniline-like and tertiary C-N stretching
1060-1040 --C-N secondary amine C-N stretching
Figure 10: Infrared vibration spectra of the as-received, --OH and --NH2 terminated FNDs with diameters of 30, 40, 50, 70, 90 and 140nm
Appendix D XPS supporting data
The XPS spectra of C and N elements for ND10 and ND90 particles are shown in Fig. 11. The N spectrum of the as-received ND10 sample shows the occurrence of absorbed nitrogen at around 401 eV, while amino-terminated samples show amide groups with an XPS peak at 400 eV which dominates the N related XPS spectra.
Figure 11: The XPS spectra of C and N elements in ND10 and ND90 FNDs are shown for as-received and amino terminated cases.
Appendix E Additional supporting data for the PL spectrum analysis
NV(--) emission fraction was determined for the as-received FNDs as a function of the applied laser power (Fig. 12). The results for the washed, --OH terminated, and --NH2 terminated FNDs are also shown for an enhanced number of laser powers in Fig. 13.
Figure 12: NV(--) emission fraction for the as-received FNDs as a function of the applied laser power for different FND sizes.
Figure 13: Photoluminescence spectrum analysis of FNDs. NV(--) emission fraction for the (a) washed, (b) --OH terminated, and (c) --NH2 terminated FNDs as a function of the applied laser intensity for different FND sizes.